OpenStudy (anonymous):

in a drawing if solar systems the scale is 1mm=500km for a planet with a diameter of 6,000 km what should be the diameter of the drawing of the planet

jhonyy9 (jhonyy9):

will be 6000/500 =12 mm

jhonyy9 (jhonyy9):

@Kito19 ok ?

jhonyy9 (jhonyy9):

or for more understandably us the rule of 3 simple so 1 mm for 500 km x mm for 6000 km -------------------- x = 6000 *1 / 500 = 6000/500 = 12 x = 12 mm hope this is understandably now sure right easy

OpenStudy (anonymous):

Thank you for explain it

jhonyy9 (jhonyy9):

np my pleasure